1. CVP Analysis; Break-even point, margin of safety: Davies’ Violins, Ltd, produces and sells a single product, violins, whose selling price is $325.00 per unit and whose variable cost is $98.00 per unit. The company's fixed expense is $47,300 per month. The current volume of sales is 350 violins per month.

  1. Determine the monthly total contribution margin at the current volume of sales.
  2. Determine the monthly net income (loss) at the current volume of sales.
  3. Determine the monthly break-even point:
  1. In units (round your answer up to the nearest whole unit)
  2. In sales dollars (round your answer to the nearest cent (i.e. 2 decimal places)
  1. What is the company’s margin of safety:
    1. In units
    2. In sales dollars
Answer #1

Total Contribution margin

= (Sales price - Variable costs) * units sold

= (325-98)*350

= 79,450

Net income = Total Contribution margin - Fixed costs

= 79,450 - 47,300

= 32,150

Break even point

In units = Fixed cost /Contribution margin per unit

= 47,300/(325-98)

= 208 units

In sales dollars = Fixed cost/Contribution margin ratio

Contribution margin ratio = 227/325 = 69.85%

= 47,300/69.85%

= 67,716.54

Margin of Safety

In units = Sales - Breakeven units

= 350 - 208

= 142 units

In sales dollars = 142*325

= 46,150